If the geometric mean of 1/4 and 4 is x+3,find the possible values of x.
Answer by richwmiller(17219)

1/4, x+3,4
1/4 = t
4 = 1/4 * r^(2)
16=r^2
r=4
x+3 = 1
x=-2
1/4,1, 4